OpenStudy (anonymous):

ok im lost i know its probally simple but please break it down for me. darla can enter the payroll for xyz company in 12 hours it takes rosie 6 hours to enter the same payroll how long will it take if they worjk together

OpenStudy (anonymous):

\[\frac{ 1 }{ \frac{ 1 }{ x }+ \frac{ 1 }{ y }}\]

OpenStudy (anonymous):

so we can let x = darla and y = rosie. what we are looking for here is there rate of work, or.. how many jobs can they complete in an hour.

OpenStudy (anonymous):

so if darla can do the 1 job in 12 hours.. how many jobs can she do in an hour?

OpenStudy (anonymous):

\[\frac{ 1 }{ 12}\] hours... right?

OpenStudy (anonymous):

so if darla can do 1/12 jobs per hour.. how many can rosie do?

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }\]

OpenStudy (anonymous):

so darla does 1/12 jobs per hour and rosie does 1/6 jobs per hour, when they work together they can complete \[\frac{ 1 }{ 12 }+\frac{ 1 }{ 6 }\] jobs per hour.

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }+\frac{ 1 }{ 12 }=\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

so together, rosie and darla can complete 1/4 jobs per hour.

OpenStudy (anonymous):

so the time it takes for them to complete the one job is given by the inverse of 1/12+1/6

OpenStudy (anonymous):

wheres does the 1/4 come from

OpenStudy (anonymous):

find a common denominator and add right? 2/12 + 1/12 = 3/12 = 1/4

OpenStudy (anonymous):

so the inverse of the rate of work is \[\frac{ 1 }{ \frac{ 1 }{ 4 } }\]

OpenStudy (anonymous):

do you know how to solve that last division problem?

OpenStudy (anonymous):

\[1\div \frac{ 1 }{ 4 } = \frac{ 1 }{ \frac{ 1 }{ 4 } }\times \frac{ 4 }{ 4 }=\frac{ 4 }{ \frac{ 4 }{ 4 } }= \frac{ 4 }{ 1 }=4\]

OpenStudy (anonymous):

good luck.

OpenStudy (anonymous):

so 4 hours

OpenStudy (anonymous):

yes, 4 hours for them to complete the job together.